find f(x) = x2 - 2x -2 find f' (x) using the long method


1
Expert's answer
2021-11-01T08:48:24-0400

Let "f(x) = x^2 - 2x -2". Let us find "f' (x)" using the long method.


"f'(x)=\lim\limits_{\Delta x\to 0}\frac{f(x+\Delta x)-f(x)}{\Delta x}
\\=\lim\limits_{\Delta x\to 0}\frac{(x+\Delta x)^2 - 2(x+\Delta x) -2-(x^2 - 2x -2)}{\Delta x}
\\=\lim\limits_{\Delta x\to 0}\frac{x^2+2x\Delta x+(\Delta x)^2 - 2x-2\Delta x -2-x^2 +2x +2}{\Delta x}
\\=\lim\limits_{\Delta x\to 0}\frac{2x\Delta x+(\Delta x)^2 -2\Delta x }{\Delta x}
\\=\lim\limits_{\Delta x\to 0}(2x+\Delta x -2 )\\=2x-2."